find the coordinates of the point equidistant from three given points A(5,1)B(-3,-7)andC(7,-1) 

Grade:10

2 Answers

Vinay Kumar
6 years ago
Consider a point D(x,y) which is equal distant from the points .As the distance is equal so ,AD=BD=CD.BY using distance formula between two points the value of "x"and"y" can be calculated....i.e.(x,y)=(2,-4)
safe india by sonu yadav
6 years ago
/centoid of triangle abc is equidistant from all the pin ts let it be D(x,y)
x=x1+x2+x3/3 =5+7-3/3=9/3=3
similarly y=y1+y2+y3/3=1-7-1/3=-7/3
D=(3,-7/3)
approve my answer